OpenStudy (anonymous):

can someone help me? Automobile tire pressure increases about 1psi for each 10F increase in air temperature. At an air temperature of 55F a tires pressure is 30psi. Write an equation that models the tires pressure as a function of air temperature.

OpenStudy (anonymous):

please can you people help me?

OpenStudy (anonymous):

for each 10F increase in air pressure you have a 1 psi increase, so for each 1 increase in air pressure you have a \(\frac{1}{10}=0.1\) increase in psi if you are going to write a linear equation we need the slope, and your slope is therefore \(0.1\)

OpenStudy (anonymous):

ooh i think i get that but how would we find the y intercept?

OpenStudy (anonymous):

at a pressure of 55 you get 30 psi, which is like saying the point \((55,30)\) is on the graph of your line

OpenStudy (anonymous):

ooh so then we use the formula y-y1=m(x-x1) to get the y intercept right?

OpenStudy (anonymous):

so you can use the "point-slope" formula to find the equation of your line. the point is \((55,30)\) and the slope is \(0.1\)

OpenStudy (anonymous):

yes you got it go ahead and finish, i can check if you like

OpenStudy (anonymous):

ok thank you!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

y=.1x+24.5? is that right?

OpenStudy (anonymous):

that is what i get, yes. you can also do the substitution trick \(y=.1x+b\) \(30=.1\times 55+b\) \(30=5.5+b\) \(b=24.5\)

OpenStudy (anonymous):

ooh i will try that next time thank you so much!!

OpenStudy (anonymous):

yw